# August 11, 2025

### v1

@@ Abstract
In this short-length paper, we will present some math that play a central role in quantum hydrodynamics and were presented by Mita in the past. In this formulation of QM, a quantity is involved which is the diffusion flux. As we will see, this quantity plays a crucial role on the evolution of wave packets. More specifically, we will initially briefly explain some results derived in Mita’s papers, by analyzing Gaussian and soliton wave packets, while our main focus is on the role of diffusion in the evolution of super-luminal wave packets.
@@ I Introduction
$E_{K,x}^{(stationary)}=\frac{(\Delta P)^{2}}{2M}=\frac{1}{2}M\int_{\mathbb{R}}{\frac{D^{2}(x)}{\rho(x)}}dx$
The diffusion flux is an indicator of the tendency of the wavepacket to diffuse. We should note also that, in the past, there was also a quantity named osmotic velocity that was discussed by Bohm and Hiley bohm1989non in the context of a stochastic version of Bohmian mechanics. This velocity was defined as follows:
@@
The role of that velocity according to the authors was to balance diffusion in stationary state systems, where there was no probability evolution over time. This velocity, according to the authors, is responsible for ”guiding” the particle in higher probability areas. In addition, it was noted that the osmotic velocity becomes infinite at the nodes of the wave function. We have to add that we know that for every position square-integrable wave function $\Psi(x)$, we have a corresponding momentum wave function $\Psi(\rho)$ which is also square-integrable. This means that $|\Psi(\rho\rightarrow\infty)|^{2}=0$. In other words, it’s very difficult to find a particle that has extremely large velocity. And indeed, this is what we observe with the osmotic velocity. The particle has very large velocity close to the nodes of the wave function, where it is obviously very difficult to be observed. And that, of course, sounds very natural since the dwell time of a particle at a specific spatial region is inversely proportional to its velocity in that region.
Now that we are finished with the osmotic velocity, we will move on to the analysis of the evolution of wave packets. In chapter II we will briefly discuss both Gaussian and soliton wave packets and at the end, in Chapter III, we will see the role of diffusion flux in the evolution of initially extremely localized quantum systems.
II DIFFUSION FLUX IN GAUSSIAN AND SOLITON WAVE PACKETS
The role of this Chapter is not to produce new knowledge, but to introduce the diffusion flux by looking specifically at the evolution of two types of waveforms. First, we have the Gaussian wave packet:
@@ II DIFFUSION FLUX IN GAUSSIAN AND SOLITON WAVE PACKETS
$\overrightarrow{J}_{g}(x,t)=\rho_{g}(x,t)u_{0}\overrightarrow{e}_{x}+\frac{mt}{T}\overrightarrow{D}_{g}(x,t)$
The first term is related to the ”centre of mass” motion of the wave packet and the second term, which contains the diffusion flux is related to the expansion of the probability distribution. Now, when it comes to the soliton wave packet, Mita mita2021schrodinger defined a probability distribution:
@@
$\overrightarrow{F}(x,t)=\frac{2\hbar^{2}}{m\sigma_{0}}\overrightarrow{\nabla}\rho_{sol}(x,t)\sim-\overrightarrow{D}_{sol}(x,t)$
And that’s how the form of the wave packet is preserved. The greater the tendency of the wave packet to diffuse (greater $\overrightarrow{D}$) the larger the external force of opposite direction that has to be applied. And, of course, the coefficient $\mu=\frac{2\hbar^{2}}{m\sigma_{0}}$ is position independent, because otherwise it would cause deformation of the wave packet.
III DIFFUSION AND SUPER-LUMINAL WAVE PACKETS
@@
$\Psi(x,0)=\frac{1}{\sqrt{a}}\chi_{[-a/2,a/2]}(x)e^{ik_{0}x}$
According to Mita again mita2007dispersion the time-evolved state in free space will be the following:
@@
$\Psi(x,0)=\frac{1}{\sqrt{a}}\Theta(a/2-|x|)e^{ik_{0}x}=\frac{1}{\sqrt{a}}\Theta(a/2-x)\Theta(a/2+x)e^{ik_{0}x}$
@@
$\lim_{x\rightarrow a/2}\overrightarrow{D}(x,0)=(+\infty)\overrightarrow{e}_{x}$
@@
$\lim_{x\rightarrow-a/2}\overrightarrow{D}(x,0)=(-\infty)\overrightarrow{e}_{x}$
@@
IV DISCUSSIONS AND CONCLUSIONS
In this paper, we managed to give some clues to support the idea of contribution of the osmotic velocity field on quantum dynamics, even on stationary states. We also presented the mathematical and physical/hydrodynamical reason behind the infinite diffusion of strictly localized wave packets. In conclusion, if someone wants to examine the evolution of a wave-packet they’d better look at the Bohmian velocity. However, as we saw, diffusion flux also provides us with significant information relevant to the origin of quantum diffusion, which arises from the diffusion tendencies of the initial state (that are expressed by the diffusion flux).

### v5

# August 27, 2025

@@ Abstract
In this short-length paper, we will present some math that play a central role in quantum hydrodynamics and were presented by Mita in the past. In this formulation of QM, a quantity is involved which is the diffusion flux. As we will see, this quantity plays a crucial role on the evolution of wave packets. More specifically, we will initially briefly explain some results derived in Mita’s papers, by analyzing Gaussian and soliton wave packets, while our main focus is on the role of diffusion in the evolution of super-luminal wave packets. Also, a very brief discussion will be made about the presence of the osmotic velocity field in quantum dynamics.
@@ I Introduction
$E_{K,x}^{(stationary)}=\frac{(\Delta P)^{2}}{2M}=\frac{1}{2}M\int_{\mathbb{R}}{\frac{D^{2}(x)}{\rho(x)}}dx$ (3)
The diffusion flux, in general, is an indicator of the tendency of the wavepacket to diffuse.
II THE OSMOTIC VELOCITY
We should note that, in the past, there was also a quantity named osmotic velocity that was discussed by Bohm and Hiley bohm1989non in the context of a stochastic version of Bohmian mechanics. This velocity was defined as follows:
@@ II THE OSMOTIC VELOCITY
. The role of that velocity according to the authors was to balance diffusion in stationary state systems, where there was no probability evolution over time. This velocity, according to the authors, is responsible for ”guiding” the particle in higher probability areas. It was also noted that the osmotic velocity becomes infinite at the nodes of the wave function. Based on that, we’ll give our own interpretation of this quantity. To be more specific, we know that for every position square-integrable wave function $\Psi(x)$, we have a corresponding momentum wave function $\Psi(\rho)$ which is also square-integrable. This means that $|\Psi(\rho\rightarrow\infty)|^{2}=0$. In other words, it’s very difficult to find a particle that has extremely large velocity. And indeed, this is what we observe with the osmotic velocity. The particle has very large osmotic velocity close to the nodes of the wave function, where it is obviously very difficult to be observed. And that, of course, sounds very natural since the dwell time of a particle at a specific spatial region is inversely proportional to its velocity in that region. In a similar way, we can explain why the osmotic velocity approaches zero in the highest probability areas. And, of course, taking into advantage of the osmotic velocity expression, Eq. (3) gives us in stationary states that $E_{K,x}^{stationary}=\frac{1}{2}M\int_{\mathbb{R}}\rho(x)u^{2}(x)dx=\frac{1}{2}M<u^{2}(x)>$, as expected.
Now that we are finished with the osmotic velocity, we will move on to the analysis of the evolution of wave packets. In chapter III we will briefly discuss both Gaussian and soliton wave packets and at the end, in Chapter IV, we will see the role of diffusion flux in the evolution of initially extremely localized quantum systems.
III DIFFUSION FLUX IN GAUSSIAN AND SOLITON WAVE PACKETS
The role of this Chapter is to introduce the diffusion flux by looking specifically at the evolution of two types of waveforms. First, we have the Gaussian wave packet:
@@ III DIFFUSION FLUX IN GAUSSIAN AND SOLITON WAVE PACKETS
$\overrightarrow{J}_{g}(x,t)=\rho_{g}(x,t)u_{0}\overrightarrow{e}_{x}+\frac{t}{T}\overrightarrow{D}_{g}(x,t)$
The first term is related to the ”centre of mass” motion of the wave packet and the second term, which contains the diffusion flux is related to the expansion of the probability distribution. After very long time, the system approaches an equilibrium state in which we observe that the wave packet diffusivity approaches zero. Indeed, let’s choose for simplicity that $u_{0}=0$ and assume that $t>>0$ and $|x|<<\infty$. We observe that since
$\overrightarrow{D}_{g}(x,t)=\frac{2\sqrt{2}\hbar}{\sqrt{\pi}m\epsilon^{3}(t)}xe^{-2x^{2}/\epsilon^{2}(t)}\overrightarrow{e}_{x}$
we obtain:
$\overrightarrow{D}_{g}(x,t>>0)\approx\left(\frac{2\sqrt{2}\hbar}{\sqrt{\pi}ma^{3}}\frac{xT^{3}}{t^{3}}\left(1-2\frac{x^{2}T^{2}}{a^{2}t^{2}}+O\left(\frac{x^{4}}{t^{4}}\right)\right)\right)\overrightarrow{e}_{x}\Rightarrow$
$\overrightarrow{D}_{g}(x,t>>0)\approx\left(\frac{2\sqrt{2}\hbar}{\sqrt{\pi}ma^{3}}\frac{xT^{3}}{t^{3}}+O\left(\frac{x^{3}}{t^{5}}\right)\right)\overrightarrow{e}_{x}\Rightarrow$
$\overrightarrow{D}_{g}(x,t>>0)\sim\frac{x}{t^{3}}\overrightarrow{e}_{x}=s\overrightarrow{e}_{x}$
where $|s|<<1$. Now, when it comes to the soliton wave packet, Mita mita2021schrodinger defined a probability distribution:
@@
$\overrightarrow{F}(x,t)=\frac{2\hbar^{2}}{m\sigma_{0}}\overrightarrow{\nabla}\rho_{sol}(x,t)\propto-\overrightarrow{D}_{sol}(x,t)$
And that’s how the form of the wave packet is preserved. The greater the tendency of the wave packet to diffuse (greater $\overrightarrow{D}$) in a specific direction, the larger the ”effective” force of opposite direction that has to be applied. And, of course, the coefficient $\mu=\frac{2\hbar^{2}}{m\sigma_{0}}$ is position independent, because otherwise it would cause deformation of the wave packet.
We have to add that this scalar ”effective” potential we referred to is present in the very well-known NLS equation for solitons zakharov1974complete , which appears a lot in Bose-Einstein condensates:
$i\hbar\partial_{t}\Psi_{sol}(x,t)=-\frac{\hbar^{2}}{2m}\partial_{x}^{2}\Psi_{sol}(x,t)-\mu|\Psi_{sol}(x,t)|^{2}\Psi_{sol}(x,t)$
IV DIFFUSION AND SUPER-LUMINAL WAVE PACKETS
@@
$\Psi(x,0)=\frac{1}{\sqrt{a}}\chi_{[-a/2,a/2]}(x)$
According to Mita again mita2007dispersion , the time-evolved state in free space will be the following:
@@
$\Psi(x,0)=\frac{1}{\sqrt{a}}\Theta(a/2-|x|)=\frac{1}{\sqrt{a}}\Theta(a/2-x)\Theta(a/2+x)$
@@
$\lim_{x\rightarrow a/2^{-}}\overrightarrow{D}(x,0)=(+\infty)\overrightarrow{e}_{x}$
@@
$\lim_{x\rightarrow-a/2^{+}}\overrightarrow{D}(x,0)=(-\infty)\overrightarrow{e}_{x}$
@@
V DISCUSSIONS AND CONCLUSIONS
In this paper, we managed to give some clues to support the idea of contribution of the osmotic velocity field on quantum dynamics, even on stationary states where Bohm’s/current velocity plays no role. We also presented the mathematical and physical/hydrodynamical reason behind the infinite diffusion of strictly localized wave packets. So, one the one hand, if someone wants to examine the evolution of a wave-packet, they’d better look at the Bohmian current velocity, but on the other hand, as we saw, diffusion flux also provides us with significant information relevant to the origin of quantum diffusion, which arises from the diffusion tendencies of the initial state (that are expressed by the diffusion flux).